OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b^2 - 4ac? 1 = 2x^2 + 5x?

OpenStudy (anonymous):

2x^2 + 5x -1 = 0 general form a = 2, b = 5 and c = -1 now u can calculate b^2-4ac on yr own.....

jhonyy9 (jhonyy9):

b*(2) - 4ac = 25 +8 = 33